OpenStudy (anonymous):

Which triple of numbers:\[x_{1},x_{2},x_{3}\in \mathbb{Z}^+\]Is a solution to:\[\left[\begin{matrix}x _{1} & x _{2} & x _{3} & 17 \\ -x _{1} & 2x _{2} & -16x _{3} & -35 \\ 3x _{1} & 4x _{2} & -2x _{3} & 45\end{matrix}\right]\]

OpenStudy (experimentx):

if you post in simple way such that i can copy, rows separated by , and column separated by ; then i will tell you.

OpenStudy (anonymous):

I got it down to a relatively simple set of equations: x_1=23-6t x_2=-6+5t x_3=t Where would I go from here?

OpenStudy (kinggeorge):

Each of the values for \(x_1, x_2, x_3\) must be less than 17, and greater than 0. So t has to be at least 2 for \(x_2\) to be positive, and less than 4 for \(x_1\) to be positive. So try integer values of t between 2 and 4, and check if your equations are satisfied.

OpenStudy (anonymous):

Thanks!

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (experimentx):

Sorry, I was kinda bg.

OpenStudy (anonymous):

No worries! :)